The perimter of a triangle is 46 cm. The triangle is isosceles now, but if its base were lenghtened by 6cm and each leg were shortened by 2 cm, it would be equilaterial. Find the base of the original triangle.
Answer by edjones(8007)

Let an isosceles side be s and the base be b.
A) 2s+b=46
B) s-2=b+6
.
B) s=b+8
A)
2(b+8)+b=46 Substitute b+8 for s.
2b+16+b=46
3b+16=46
3b=30
b=10 Ans.
s=18
.
Check:
Sides of original triangle: 18, 18, 10
Sides of equilateral triangle: 16, 16, 16
.
Ed